OpenStudy (anonymous):

Factoring cubes/Solve by factoring: 1. w^3-2y^3 2. a^8-a^2b^6 3. 12qw^3-12q^4

OpenStudy (anonymous):

\[w ^{3}-2y ^{3} \] \[a ^{8}-a ^{2}b ^{6}\] \[12qw ^{3}-12q ^{4}\]

zepdrix (zepdrix):

\[\large (a^3-b^3)=(a-b)(a^2+ab+b^2)\] So this is the formula for difference of cubes, we'll need to use this to get through these :D

OpenStudy (anonymous):

Yes. I have the formula. Sounds good. I was able to do some of the problems but I'm having a hard time with these three

zepdrix (zepdrix):

Hmm the first one is a little annoying. Since we have that 2 there. It will make things a little ugly. Let's see if we can identify a and b.

OpenStudy (anonymous):

YES - I do not like the 2

zepdrix (zepdrix):

\[\large a=w\]\[\large 2y^3=(\sqrt[3]{2}y)^3 =b^3\]\[\large b=\sqrt[3]{2} y\] Understand what we did there? :o it was a little tricky.

zepdrix (zepdrix):

We needed to rewrite the 2 as a cube. So we could think of it as the difference of cubes.

OpenStudy (anonymous):

It is a bit confusing but OK

zepdrix (zepdrix):

sec, i'm trying to think if there is a better way we can write it :)

OpenStudy (anonymous):

What is confusing me is thinking of how to continue the formula with a radical cube

zepdrix (zepdrix):

We continue the way we normally would, it just won't look very pretty :D \[\large =(w-\sqrt[3]{2}y)(w^2+\sqrt[3]{2}wy+y^2)\]